OpenStudy (anonymous):

\(\sqrt{81m^3}\) Simplify Radical

OpenStudy (lgbasallote):

factor out into prime factors \(\sqrt{(3)(3)(3)(3)(m)(m)(m)}\) take out of the square root all that appeared twice

OpenStudy (anonymous):

confused >.<

OpenStudy (lgbasallote):

i just factored it out..81 = 9 x 9 right? and m is (m)(m)(m) \(\sqrt {(9)(9)(m)(m)(m)}\) better? now take out of the radical the numbers/letters that appeared twice..9 and m appeared twice so take out of the square root 2 of them \(9m \sqrt{m}\) do you get that method?